OpenStudy (anonymous):

if 1 brick weighs 5 pounds how if 1 brick weighs 5 pounds how @Mathematics

OpenStudy (anonymous):

much does does 12 weigh

hero (hero):

1/5 = 12/x 60 = x 60 pounds